# Pixel-Wise Recognition for Holistic Surgical Scene Understanding

### v2

@@ 5.2.2 Instrument Segmentation Experiments
Due to our new instance segmentation annotations and the data curation process, our annotations in GraSP have more precise bounding box coordinates than the PSI-AVA benchmark. In this section, we first assess the performance of multiple object detectors on these bounding box annotations, and we use these results as reference points for our experiments with instance segmentation. For this purpose, we adapt, train, and optimize multiple object detectors on our cross-validation set and present our results in Table 11. We compare Faster R-CNN (Ren et al. (2017)) as a CNN-based detector with Deformable DETR (Zhu et al. (2021)) as a transformer-based detector. In this work, we also test DINO (Zhang et al. (2023)) with its largest backbone (Swin Large (Liu et al. (2021))) as a way of including a larger and more recent state-of-the-art method in the evaluation of instrument detection performance. Our results show that transformer-based detectors consistently surpass CNN-based detectors and that the DINO detector performs better than Deformable DETR. These results are consistent with our initial results (Valderrama et al. (2022)) and with the general state-of-the-art in object detection, thus showing the reliability of our data and its applicability for instrument detection tasks.
Furthermore, we validate our segmentation annotations by training and optimizing ISINet (González et al. (2020)) as a CNN-based baseline and comparing it with our instrument segmentation baseline (called TAPIS Frame), as detailed in Table 12. We evaluate the performance of TAPIS Frame using multiple backbones, including ResNet50, for comparison with ISINet and a Swin Large (SwinL) backbone to leverage the most extensive backbone configuration. Our findings consistently demonstrate the superiority of transformer-based methods, with performance positively correlating with the number of parameters of the backbone. Notably, all instance segmentation methods exceed their detection counterparts in $mAP@0.5IoU_{box}$, with even the most minor TAPIS Frame configuration achieving comparable performance to the largest detector with less than half of the parameters and FLOPs. These results highlight our segmentation annotations’ impact on training specialized models that capture finer visual cues regarding the pixel-level shape of instruments that improve instrument localization and recognition. Additionally, as shown in Table 12, we test multiple publicly available instrument segmentation methods on our cross-validation set. For this purpose, we optimize and adapt methods designed for instrument segmentation in Endovis 2017 and 2018 by using their official public implementations; these methods include TernausNet (Shvets et al. (2018)) [^12] https://github.com/ternaus/robot-surgery-segmentation and MF-TAPNet (Jin et al. (2019)) [^13] https://github.com/keyuncheng/MF-TAPNet for semantic segmentation, and ISINet (González et al. (2020)) [^14] https://github.com/BCV-Uniandes/ISINet and QPD Mask DINO (Dhanakshirur et al. (2023)) [^15] https://github.com/AINeurosurgery/Learnable-QPD-for-maskDINO for Instance Segmentation. These methods yield a relative behavior that aligns with the current state-of-the-art in instrument segmentation and underscores the enhanced performance of our instrument segmentation model over previous approaches. Also, they consistently prove the superiority of Transformer-based architectures.
@@
Moreover, Table 12 also presents the performance of the entire TAPIS model for instrument segmentation (called TAPIS Full), which integrates the region classification head with the TAPIS Frame’s region proposals. The metrics for this experiment, including parameters and FLOPs, focus on the video feature extractor and the region classification head. Including the region classification head enhances performance in the highly stringent semantic segmentation metrics, although with a decrease in $mAP@0.5IoU_{segm}$ performance. We attribute this outcome to the head’s ability to accurately reclassify more prominent and complex regions by correlating them over time. However, it tends to misclassify smaller, independent regions that appear momentarily, resulting in higher recall regarding intersection over union but slightly reduced effectiveness in instance identification. Throughout our experimentation, we noted minimal variance in instrument classification performance related to the specific design of the time window and the region classification head. However, we observed marginally lower performance with windows covering less than 8 seconds and achieved the optimal performance using a sampling rate of 2fps. Consequently, we opted for a window design comprising 16 frames at a 2fps sampling rate.
We provide detailed results of per-class instrument segmentation performance in the Supplementary Material. On the one hand, the per-class performance is highly correlated to the frequency of each instrument type within the dataset. On the other hand, the results are significantly affected by the specific identity of each instrument and its typical presence in surgeries. For example, the MCS demonstrates excellent performance due to its distinctive shaft and tip and its frequent, clear visibility in the frontal parts of the video. Similarly, besides its unique tip, the LND often appears alone during parts of the surgery where few other instruments are present, thus underlining the importance of the visual surgical context in segmentation accuracy. On the contrary, the PF exhibits lower performance due to frequent occlusions and their appearance in distant, smaller parts of the image. Besides often being covered by other instruments or objects, the LI presents high intra-class variability. It appears in various forms throughout our data as it is a generic category for external laparoscopic instruments independent from the da Vinci Surgical System. This variability in performance further emphasizes the complex nature of instrument segmentation in surgical videos and the necessity of considering the entire visual and temporal context.
@@
Figure 11 shows examples of qualitative results of TAPIS for instrument segmentation, and we present more examples in the Supplementary Material. These results demonstrate that TAPIS achieves a very high segmentation performance in both instrument recognition capabilities and pixel-level delineation precision. For example, the results on the left part of Figure 11 demonstrate TAPIS’s ability to accurately segment complex images with up to five instrument instances with intricate, varied, and occasionally overlapping figures. The second row on the left shows an example of a correct segmentation with significantly overlapping instances. Nevertheless, our method presents some error patterns related to high occlusions and complicated instrument appearances. The results on the first three rows on the right show minor defects on moving and still instruments occluded by tissue and fluids. The last three rows on the right show examples of inaccurate predictions due to a high degree of overlapping between instances and a problematic view of specific instruments, which makes them challenging to segment accurately, even for the naked eye. However, we highlight that TAPIS achieves a very high segmentation recall as it can propose a segmentation mask to cover most instances, failing primarily in tiny segments. These insights prove our proposal’s strength but show room for improvement in segmentation.
@@ 5.2.3 Atomic Action Detection
We first evaluate the impact of using different region proposal networks for the action detection tasks. Table 13 contains the results of testing TAPIS’s atomic action detection performance and our proposed region classification head with region features provided by multiple transformer-based architectures. We also report the number of parameters and floating point operations of those region proposal architectures. The results demonstrate that performance in atomic action detection increases proportionally to instrument detection or segmentation performance. This relation occurs since higher-quality region proposals provide a broader instrument location recall and better recognition insights in their features, leading to increased action detection and recognition capabilities. Additionally, we observe that using TAPIS Frame’s features with a ResNet50 backbone achieves a comparable performance to DINO’s features with a SwinL backbone. These results highlight the impact of segmentation approaches as they provide finer-grained features with detailed cues of instruments’ shapes and poses that improve TAPIS’s understanding of the actions that those instruments perform. On top of that, for all the experiments, we used the pretrained weights in phase recognition as we observed that these learned models provided better results for atomic action detection.
@@
SwinL 868 216 34.04 ±$\pm 2.40$
Moreover, we assess our proposed region classification head compared with TAPIR’s design (Valderrama et al. (2022)). TAPIR’s box classification head performed a mean pooling of spatio-temporal features and concatenated the result with the region embeddings for a final linear classification. We now evaluate the efficacy of employing a cross-attention layer that correlates the region features with the spatio-temporal features of the video feature extractor. For a fair comparison, given the higher parameters and floating point operations count of the cross-attention layer, we also test using a multi-layer perceptron (MLP) similar in design to the one presented in MATIS (Ayobi et al. (2023)), which consists of multiple linear layers before and after the pooling operation. Table 14 shows the results of using TAPIR’s original design (Time Pooling), MATIS’s design (Time MLP), and our proposed cross-attention layer (Cross-Attention). The results show that, for any region proposal architecture, the performance scales with the number of parameters as the Time MLP outperforms the Time Pooling, thus showing consistency with the results reported in Ayobi et al. (2023). More importantly, the cross-attention layer significantly outperforms both Time MLP and Time Pooling designs despite having slightly lower Parameters and FLOPs counts than the Time MLP. This improvement demonstrates the effectiveness of our proposed design, where the cross-attention layer allows region embeddings to attend to all the features extracted by the video feature extractor, hence avoiding the heavy loss of information caused by pooling the rich sequence of temporal embeddings. Therefore, our proposed region classification head performs better since it classifies features enriched by observing the complete temporal information provided by the video feature extractor.
Table 14: Ablation results on the design of the region classification head in the cross-validation set. We compare the performance in atomic action detection using the Time Pooling design of Valderrama et al. (2022), the Time MLP from Ayobi et al. (2023), and our new Cross-Attention layer design. The best results are shown in bold.
@@
TAPIS Frame (SwinL) 32.48 ±$\pm 1.23$ 33.43 ±$\pm 1.38$ 34.04 ±$\pm 2.40$
@@
Finally, we conduct a series of design experiments to achieve the optimal time window size for atomic action detection, and we present our results in Table 15. Our observations reveal that our model performs best when the window size ranges between 16 and 32 frames, and we note a significant drop in performance when using a smaller window size. We attribute this decline to the model’s reduced temporal field, compromising its ability to capture atomic actions that extend over longer periods, such as tissue cauterization or cutting. Conversely, larger time windows also resulted in a slight decrease in performance. We hypothesize that this decline is due to the model processing noise from irrelevant distant frames, which complicates the correlation between region features and temporal embeddings. These results underscore the need for a more localized and centered temporal understanding of atomic action recognition.
Additionally, our experiments showed that decreasing the temporal window’s sampling rate per second (increasing the temporal stride) significantly decreases performance, hence validating our hypothesis about the temporal granularity of the atomic actions detection task. As their name indicates, atomic actions develop in extremely short periods, and they are a highly specific and granular form of modeling surgical activities. During the annotation process, we observed numerous instances where atomic actions unfolded in as few as four consecutive video frames. This observation indicates that the model requires a finer temporal sampling to fully identify actions over time. Consequently, the model performs best with the highest possible sampling rate, allowing for a more detailed and accurate temporal analysis of the surgical actions. This insight is crucial for enhancing the precision and effectiveness of atomic action detection in surgical videos.
Table 15: Temporal window design experiments for atomic action detection. The number of frames denotes the temporal window size centered on the frame of interest. Temporal stride is the gap between consecutive sampled frames in seconds.
Window Size Temporal Stride (seconds) $mAP@0.5IoU_{box}$ (%)
8 0.033 32.74 ±$\pm 2.12$
8 0.066 $3$$\pm 1.74$
16 0.033 34.04 ±$\pm 2.40$
16 0.066 33.37 ±$\pm 1.59$
16 0.5 30.32 ±$\pm 1.60$
16 1 28.07 ±$\pm 2.78$
32 0.033 33.66 ±$\pm 1.10$
32 0.066 33.58 ±$\pm 0.42$
@@
We provide a table with the per-class results of the atomic action detection task in the Supplementary Material. The performance distribution is once again related to the frequency of each action’s occurrence within the dataset since actions like travel, hold, and still demonstrate the highest performance metrics. However, the complexity and variability of the actions also significantly influence their detectability. For example, actions like cauterize and suction record good $AP$ values, as they are specific actions associated with particular instruments. Also, contextual cues provided by the video feature extractor, such as the appearance of smoke during cauterization or the movement of blood during suction, further aid in recognizing these actions. Conversely, while seemingly straightforward to identify, actions like pull and push often get confused with each other or with similar actions, leading to relatively lower performance. Moreover, abstract action categories such as release or open something can present in many ways within the surgical environment, making it challenging to consistently classify these actions accurately, as reflected in their low-performance metrics. This aspect of our results highlights the intricate nature of action detection in surgical videos and the need to precisely understand subtle and significant differences and overlaps in the motion patterns.
@@ 5.3 Discussion
The outstanding performance of TAPIS in different tasks and the consistency observed throughout our experiments in alternative frameworks and the GraSP dataset strongly emphasizes the efficacy of our proposed benchmark and methods. Our experimentation effectively showcases how our approach leverages the diverse granularity requirements of each task within the proposed multi-level hierarchy of surgical procedures. For instance, at the highest level of our task pyramid, phases recognition needed global and sparse time windows (64s at 0.25fps) to capture long-term procedural information effectively. Moving down the hierarchy, steps recognition achieved optimal performance with finer sampled time windows (16s at 1fps). Instrument recognition excelled with the most granular spatial recognition (pixel-wise) coupled with a relatively short and dense temporal field (8s at 2fps). Finally, at the base of the pyramid, atomic action detection, our lowest-level task, saw significant performance enhancements when applying the finest-grained approaches in time (0.53s at 30fps) and space.
Our results also confirm the hypothesized interrelationships among the tasks. We observed that using pretrained weights on the Phase recognition task could enhance performance in shorter-term tasks. Similarly, improving the recall in instrument recognition and employing shape-level recognition positively impacted the effectiveness of atomic action detection. This interconnected approach reinforces our understanding of the hierarchical nature of surgical scene analysis. Finally, TAPIS’s superiority in the GraSP dataset and alternative benchmarks highlights our proposed method’s robustness and ability to efficiently leverage multi-level annotations in endoscopic vision datasets. Although TAPIS achieves state-of-the-art performance across ours and alternative benchmarks, our results exhibit considerable room for improvement in all the proposed tasks of our framework. Still, our methodologies set a strong precedent and significantly push the envelope for future research in our benchmark to achieve a more holistic and multi-granular surgical scene understanding.
@@ 6 Conclusions
We have introduced a novel dataset and methodological framework to formulate surgical scene understanding into a holistic comprehension. Our approach models surgical procedures as a hierarchy of complementary long-term and short-term tasks, implicitly requiring multiple levels of visual understanding. By curating and enlarging our previously proposed benchmark for holistic surgical scene understanding and incorporating the surgical instrument instance segmentation task, we have presented the first Endoscopic Vision dataset that provides four distinct levels of visual understanding, each with the highest annotation granularity. Through detailed statistics of our dataset, we showed that our benchmark poses multiple challenges that are highly pertinent to real-life applications, effectively capturing the intricacies of robot-assisted surgical procedures. Additionally, we have established an explicit and reliable training, validation, and testing distribution to promote future reproducibility and comparability in our framework.
Furthermore, we developed the TAPIS model to exploit the provided annotations of our GraSP benchmark. Our results demonstrated the effectiveness of instrument segmentation annotations, allowing for better instrument recognition capabilities and providing helpful shape-level cues for better atomic action detection. Similarly, our extensive design experimentation with TAPIS highlighted each task’s particular granularity requirements and the intrinsic complementarity among different tasks. Higher-level tasks achieved the best performance with sparser time windows, and lower-level tasks outperformed with finer spatial recognition and shorter-term temporal processing. Our method significantly performed all alternative baselines and previously proposed methods, thus underscoring the superiority of transformer-based approaches and setting a new state-of-the-art in holistic surgical scene understanding. We validated our benchmark and method’s reliability by showing that TAPIS achieves a consistent relative performance in public benchmarks. Our work presents a significant advancement in the Endoscopic Vision field, aiming to promote a multi-granular and multi-level understanding of surgical procedures. This research is a stepping stone in holistic surgical scene understanding and image-guided surgical robot intelligence.

### v3

@@ 5.2.2 Instrument Segmentation Experiments
We trained and optimized multiple object detectors on our cross-validation set to assess the quality of our enhanced bounding box annotations and establish reference points for instance segmentation. The results in Table 11 compare Faster R-CNN ([71]) as a CNN-based detector with Deformable DETR ([94]) and DINO ([91]) with its largest backbone (Swin Large ([55])) as transformer-based detectors. Transformer detectors consistently surpass CNN-based detectors, with DINO achieving the best performance. These results are consistent with our initial results ([84]) and with the general state-of-the-art in object detection, thus showing the reliability of our data and its applicability for instrument detection tasks.
We validate our segmentation benchmark by comparing various configurations of TAPIS with publicly available methods for instrument segmentation ([78, 41, 32, 18]). As shown in Table 10, we assess the performance of our sole instrument segmentation baseline (referred to as TAPIS Frame) using different backbones, including ResNet50 for a fair comparison with alternative methods and Swin Large (SwinL). Transformer-based methods consistently outperform CNN-based models, improving performance as the backbone complexity increases. Notably, all TAPIS Frame configurations surpass their detection counterparts in $mAP@0.5IoU_{box}$, with even the smallest TAPIS Frame setup achieving performance comparable to the largest detection models while requiring significantly fewer parameters and FLOPs. These results demonstrate how our segmentation annotations support training pixel-level models to capture finer visual details that enhance instrument localization and recognition. Moreover, the relative performance of the compared methods aligns with trends in the current state of the art, with TAPIS achieving the best performance.
@@
Table 10 also presents the results of the entire TAPIS model for instrument segmentation (called TAPIS Full), which integrates region classification head with TAPIS Frame’s region proposals. Including the region classification head enhances the stringent semantic segmentation metrics but decreases the $mAP@0.5IoU_{segm}$. This trade-off arises from the head’s ability to reclassify more prominent regions over time while occasionally misclassifying smaller, momentary regions. Throughout our experimentation, we noted minimal variance in instrument classification performance related to the specific design of the region classification head design.
We detail the per-class instrument segmentation results in Tables 7 and 8 of the Supplementary Material. Performance strongly correlates with instrument frequency in the dataset and their visual characteristics. Instruments like the BF, MCS, and LND demonstrate high accuracy due to their distinctive shafts and tips and their frequent, unobstructed foreground visibility. In contrast, the PF exhibits lower performance as it is frequently occluded and often appears as a smaller or distant object within the frame. At the same time, the LI faces challenges from high intra-class variability as it is a generic category for external laparoscopic instruments. These variations show the complexity of surgical instrument segmentation and the importance of leveraging both visual and temporal context to achieve accurate results.
@@
Figure 11 shows examples of qualitative results of TAPIS for instrument segmentation, and we present more examples in Figures 18, 19 and 20 of the Supplementary Material. The results on the left part of Figure 11 demonstrate TAPIS’s ability to accurately segment complex images with up to five instrument instances with intricate, varied, and occasionally overlapping figures. Nevertheless, the results on the first three rows on the right show minor defects on instruments occluded by tissue and fluids. The last three rows on the right show inaccurate predictions due to a high degree of overlapping between instances and a complex views of specific instruments. However, we highlight that TAPIS achieves a very high segmentation recall as it can propose a segmentation mask to cover most instances, failing primarily in tiny segments.
@@ 5.2.3 Atomic Action Detection
We first evaluate the impact of using different region proposal networks for the action detection tasks. The results in Table 12 demonstrate that atomic action detection performance increases with instrument detection or segmentation performance, as higher-quality region proposals increase instrument location recall and provide richer semantic region features for action recognition. Additionally, using TAPIS Frame’s segmentation features with a ResNet50 backbone achieves comparable performance to DINO’s detection features with a SwinL backbone, demonstrating the value of segmentation approaches in providing more detailed features with instrument shape cues that improve atomic action comprehension.
@@
SwinL 868 216 35.46 ±$\pm 2.40$
We assess the effectiveness of our cross-attention-based region classification head against the original designs of TAPIR ([84]) and MATIS ([5]), with results summarized in Table 13. TAPIR’s classification head pools spatio-temporal features and concatenates them with region embeddings (Time Pooling in Table 13), while MATIS introduces a Time MLP with additional linear layers around TAPIR’s pooling operation (Time MLP in Table 13). Our findings reveal that performance scales with increasing parameters, as Time MLP outperforms Time Pooling, consistent with previous results in [5]. More importantly, our cross-attention layer significantly outperforms both designs despite having slightly fewer parameters and FLOPs than Time MLP. This enhancement demonstrates the superiority of our approach, as the cross-attention mechanism effectively leverages the full spatio-temporal context from the video feature extractor, avoiding the information loss associated with pooling operations and leading to more robust classification performance.
Table 13: Ablation results on the design of the region classification head in the cross-validation set. We compare the performance in atomic action detection using the Time Pooling design of [84], the Time MLP from [5], and our new Cross-Attention layer design. The best results are shown in bold.
@@
TAPIS Frame (SwinL) 32.48 ±$\pm 1.23$ 33.43 ±$\pm 1.38$ 35.46 ±$\pm 2.40$
@@
We provide the per-class results of the atomic action detection task in Table 9 of the Supplementary Material. The performance once again correlates with the frequency of each action, where actions like travel, hold, and still demonstrate the highest performance metrics. However, performance is also influenced by the complexity and visual variability of actions. Specific actions such as cauterize and suction record high AP values due to their association with distinct instruments and contextual cues from the video feature extractor, such as smoke or blood movement, that enhance their recognition. Conversely, actions like pull and push often get interchanged, leading to lower performance. Similarly, abstract actions such as release or open something can present in multiple ways, making it challenging to classify them consistently. These results underscore the complexity of surgical atomic action detection and the importance of capturing nuanced motion patterns to differentiate similar activities.
Table 14: Ablation results on atomic action fine-tunning. We evaluate the effect of Long-Term tasks pretraining and using pre-classified region embeddings (Class Regions) for atomic action detection.
Model Class Regions Long-Term Pretrain $mAP@0.5IoU_{box}$
Regions MLP 20.96 ±$\pm 0.53$
22.28 ±$\pm 0.14$
TAPIS 33.52 ±$\pm 1.05$
34.06 ±$\pm 1.95$
35.46 ±$\pm 0.34$
Table 14 evaluates the impact of instrument type segmentation and phase and step recognition signals on atomic action detection. To analyze the role of instrument recognition, we trained TAPIS using unclassified instrument regions obtained from our best segmentation baseline trained to generate generic instrument masks without a classification loss. To assess the impact of independent region features, we trained an MLP directly on region features to predict atomic actions without considering temporal processing (Regions MLP in Table 14). Results indicate that region features alone achieve notable performance as they capture rich instrument shape and pose information, facilitating action recognition. Similarly, both the Regions MLP and TAPIS demonstrated improved performance when classified regions were used, proving the importance of instrument type information for atomic action detection. We also compare TAPIS’s performance without pretraining the video feature extractor on phase and step recognition and only using MViT’s pretraining in kinetics400. The results confirm that long-term task pretraining enhances atomic action detection, highlighting the complementary nature of GraSP’s tasks and TAPIS’s capability to leverage cross-task correlations for improved performance on individual tasks.
@@ 5.3 Discussion
TAPIS’s strong performance in GraSP and its generalization on alternative frameworks demonstrates its versatility and suitability for multi-granular surgical modeling. The limited adaptability of SOTA models to varying granularities underscores the need for versatile architectures that generalize across task formulations and maintain competitive performance. TAPIS also proved versatility regarding the choice of video feature extractor or region proposal model, delivering competitive results across long- and short-term tasks with different configurations.
Moreover, our results confirm the hypothesized interrelationships among tasks in surgical workflow analysis. First, training long-term tasks simultaneously enhanced their per-task performance, enabling a single model to tackle both tasks while being trained once. In turn, this long-term task pretraining boosted short-term task performance. Similarly, improved instrument recall, using shape-level features and incorporating instrument class in region embeddings, positively impacted atomic action detection. These findings highlight TAPIS’s ability to leverage multi-task signals to enhance individual task performance. Additionally, our results underscore the importance of studying endoscopic vision and surgical workflow analysis tasks as interconnected components rather than isolated objectives.
Finally, TAPIS’s offline nature limits its application in real-time scenarios, confining its immediate utility to retrospective tasks such as data analysis, annotation, or surgical training. However, rather than prioritizing immediate assistive functionality, TAPIS and GraSP introduce groundbreaking approaches to holistic surgical scene understanding by exploiting the inherent complementarity among surgical workflow analysis tasks. This shift transcends the conventional focus on isolated tasks, establishing a foundational framework for future robotic systems capable of complex surgical assistance and cognition through a multi-level understanding of workflows and a fine-grained recognition of surgical agents and actions.
@@ 6 Conclusions
We introduced a novel dataset and framework for holistic surgical scene understanding with a hierarchy of complementary long-term and short-term tasks with varying levels of temporal and spatial detail. By curating and expanding our previous benchmark and including surgical instrument instance segmentation, we introduced the first Endoscopic Vision dataset with four levels of visual understanding within their highest annotation granularity. Through detailed dataset statistics, we highlighted the challenges presented by GraSP and its representation of real-world complexities in robot-assisted surgical procedures. We also established training, validation, and testing splits to ensure consistent benchmarking.
We developed the TAPIS model to leverage GraSP’s multi-granular annotations. TAPIS surpassed alternative baselines and SOTA models in multiple benchmarks, proving its versatility to varying task granularities and formulations. Our instrument segmentation annotations significantly improved instrument recognition and provided shape-level cues that enhanced atomic action detection. Our experiments also revealed the complementarity among tasks as TAPIS enhanced per-task performance by leveraging multi-task signals.
TAPIS’s transformer-based design consistently outperformed alternative baselines, proving the potential of transformer-based approaches and establishing new SOTA results in our benchmark. Moreover, its consistent performance across public benchmarks validates the reliability of our methodology and dataset. Hence, this work represents a significant step forward in Endoscopic Vision, promoting a multi-granular understanding of surgical procedures and laying the groundwork for holistic surgical scene understanding and image-guided surgical robot intelligence.